OpenStudy (anonymous):

help me plzz solve for x 5/(2x+3)=4/x

OpenStudy (anonymous):

\[\frac{5}{2x+3}=\frac{4}{x}\] \[5x=4(2x+3)\]

OpenStudy (kira_yamato):

5/(2x+3)=4/x 5x = 4(2x+3) [Cross Multiply] Expand and solve for x

OpenStudy (texaschic101):

exactly....cross multiply because this is a proportion

OpenStudy (anonymous):

idk what it is that is why i asket it

OpenStudy (texaschic101):

5/2x + 3 = 4/x cross multiply (5)(x) = 4(2x + 3) -- distribute through the parenthesis 5x = 8x + 12 -- subtract 8x from both sides 5x - 8x = 8x - 8x + 12 -- combine like terms -3x = 12 -- divide both sides by -3 x = -4 check.. 5(x) = 4(2x + 3) 5(-4) = 4(2(-4) + 3) -20 = 4(-8 + 3) -20 = 4(-5) -20 = -20 (correct) x = -4 any questions ?

OpenStudy (texaschic101):

do you have any questions at all ?

OpenStudy (anonymous):

no ty

OpenStudy (texaschic101):

your welcome :)